OpenStudy (anonymous):

the sum of three consecutive even integers is -54 i found -18, and -16. But i can't find -20

OpenStudy (whpalmer4):

sum of 3 consecutive even integers is -54 What do you mean that you can't find -20? \[(-20) + (-18) + (-16) = -54\]right? Here's how I would solve this: let \(x\) be the first integer. then the next even integer is \(x+2\) and then the 3rd one is \(x+4\) (all assuming that the first integer was even!) \[x + (x+2) + (x+4) = -54\] \[x + x + x + 2 + 4 = -54\]\[3x + 6 = -54\]\[3x = -54 -6\]\[3x = -60\]\[x = -20\]remember, \(x=-20\) was the first integer, so \(x+2 = -20+2 = -18\) is the second integer, and \(x+4 = -20+4 = -16\) is the third integer.

OpenStudy (anonymous):

thanks, i appreciate it, i'm very tired and my brain wasn't working so I came here to get some clarity. Thank you again!

OpenStudy (anonymous):

isn't working*...